- Press Release
- August 19, 2022
NASA ISS On-Orbit Status 25 May 2008
All ISS systems continue to function nominally, except those noted previously or below. Sunday – crew off duty. Ahead: Week 6 of Increment 17.
The CDR conducted the routine daily servicing of the SOZh system (Environment Control & Life Support System, ECLSS) in the SM, including the weekly collection of the toilet flush (SP) counter and water supply (SVO) readings for calldown to TsUP. [Regular daily SOZh maintenance consists, among else, of checking the ASU toilet facilities, replacement of the KTO & KBO solid waste containers and replacement of EDV-SV waste water and EDV-U urine containers.]
Sergey also gathered weekly data on Total Operating Time & “On” durations of the Russian POTOK-150MK (150 micron) air filter unit of the SM’s SOGS air revitalization subsystem for reporting to TsUP.
At ~10:30am EDT, FE-2 Garrett Reisman had his weekly PFC (Private Family Conference) via S-band/audio and Ku-band/MS-NetMeeting application (which displays the uplinked ground video on an SSC laptop),
The crew conducted their regular 2.5-hr. physical workout program (about half of which is used for setup & post-exercise personal hygiene) on the TVIS treadmill (CDR, FE-1, FE-2) and resistive exercise device (CDR, FE-1, FE-2).
Working off his discretionary “time permitting” task list, Oleg had his fourth run with the Russian DZZ-2 "Diatomeya" ocean observations program, using the NIKON F-5 digital still camera with 80-200 mm lens and the HDV (high-definition) video camcorder from SM windows 7 & 8 to identify areas with intensive bioproduction confined to the largest upwelling of oceanic floors. [Uplinked target zones were in the Indian Ocean (Agulhas underwater plateau – water area of the Central Depression) and the South Atlantic (Falklands plateau water area – costal area of Gabon).]
For Sergey Volkov, the Russian voluntary task list suggested another ECON KPT-3 session, making observations and taking aerial photography of environmental conditions for Russia’s Environmental Safety Agency (ECON) using the Nikon D2X digital camera with SIGMA 300-800mm telephoto lens. [Today’s target zone: the contaminated coastal area in the Persian Gulf.]
More unpacking of US cargo from Progress 29P was on Reisman’s “job jar” task list for today.
RS ASU Update: While using the SM toilet facility (ASU) on Wednesday last week, the crew heard a loud noise and later received the separation LED (Light Emitting Diode) indication. They removed & replaced the separator which is a part of the whole pump package, but the system was still exhibiting degraded suction. The decision was made today to have the crew R&R the MNR (Micropump Separator) with the last spare they had onboard. Preliminary indications are that the R&R of the MNR has fixed the problem, although specialists continue to monitor the situation. Preparations are in work to fly a replacement MNR on 1J if requested.
SKV-2 Troubleshooting Update: A small amount of coolant was discharged in the morning of 5/23 while the crew attempted to re-fill the SM SKV-2 air conditioner with Freon-218 (Khladon). The crew reported that a valve did not fit correctly resulting in the discharge. The crew quickly realized this and plugged the fitting. Ground specialists had the crew stand down on this activity for today and were assessing the leaky fitting. The SKV-2 coolant was delivered via 29P, to recharge the unit after it experienced a leak on 4/29.
Weekly Science Update (Expedition Seventeen — Week 5)
ALTCRISS (Alteino Long Term monitoring of Cosmic Rays on the ISS): Measurements continue in FGB module.
ANITA (Analyzing Interferometer for Ambient Air): Continuing.
BCAT-3 (Binary Colloidal Alloy Test 3): Reserve.
CCISS (Cardiovascular & Cerebrovascular Control on Return from ISS): Reserve.
CFE (Capillary Flow Experiment): Reserve.
CW/CR (Cell Wall/Resist Wall) in EMCS (European Modular Cultivation System): Harvesting of B3 and B4 was completed on 05/23. “Thank you for your diligent work on CW/RW and also for the last-minute gathering of EMCS hardware for return on 1J. The final activity remaining for CW/RW is the relocation of the KFTs from the +2C dewar to a -95C dewar in MELFI.
CSI-2/CGBA (CGBA Science Insert #2/Commercial Generic Bioprocessing Apparatus): In progress.
CGBA-2 (Commercial Generic Bioprocessing Apparatus 2): Complete.
CSLM-2 (Coarsening in Solid-Liquid Mixtures 2): Complete.
EarthKAM (Earth Knowledge Acquired by Middle School Students): Complete.
ELITE-S2 (Elaboratore Immagini Televisive – Space 2): Planned.
EPO (Educational Payload Operations): Reserve.
ETD (Eye Tracking Device): Completed.
EuTEF (European Technology Exposure Facility): The experiment platform continues to operate nominally, although the platform is rather cold due to unfavourable beta angles. A DHPU (Data Handling Processing Unit) software patch is currently developed to fix the link error issues encountered with DEBIE-2 and FIPEX instruments. – DEBIE-2: Link error still under investigation; – DOSTEL: On-going science acquisition; – EuTEMP: Currently inactive as planned; – EVC: Inactive; – EXPOSE: On-going science acquisition; – FIPEX: A new script was uplinked on 05/19, but it stopped too quickly on 05/20. A modified script was then uplinked on 05/21, started on 05/22 and is now scheduled to run up to 05/27; – MEDET: On 05/18, the instrument had to be switched off because it reached the low temperature limit. Awaiting more favourable thermal conditions to restart science acquisition; – PLEGPAY: Instrument is powered on, but not in science acquisition mode; – TRIBOLAB: The Pin On Disk run #3 (POD#3) was restarted nominally on 05/17 and has been running nominally since then.
FSL (Fluid Science Laboratory): FSL MIL Bus cable repair was successfully performed on 5/5. After FSL Rack Activation from ground on 5/8, the ground confirmed that the FSL cables repair activities were successful. Further troubleshooting activities will be required with CEM-U (Upper) Optical Module.
GEOFLOW: Further troubleshooting activities with CEM-U (Upper) Optical Module are required. A new lamp set is manifested on STS-124/1J. GEOFLOW start is pending further FSL troubleshooting, and remaining commissioning activities (check of FSL optical modes). This will not occur prior to 1J flight.
IMMUNO (Neuroendocrine & Immune Responses in Humans During & After Long Term Stay at ISS): Complete.
InSPACE-2 (Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions 2): In progress.
Integrated Immune: In progress.
KUBIK-FM1/ KUBIK-FM2 Centrifuge/Incubators: Completed.
LOCAD-PTS (Lab-on-a-Chip Application Development-Portable Test System): Complete.
MISSE (Materials ISS Experiment): Ongoing.
MTR-2 (Russian radiation measurements): Passive dosimeters measurements in DC1 “Pirs”.
MSG-SAME (Microgravity Science Glovebox): Complete.
NOA-2 (Nitric Oxide Analyzer): Planned.
NUTRITION w/REPOSITORY: “Garrett, thanks for your continued attention to detail in completing your FD60 Nutrition/Repository session. This was your final in-flight sample collection. We appreciate your efforts in keeping to the science/timeline constraints during blood draw procedures. Thanks for your participation!”
PMDIS (Perceptual Motor Deficits in Space): Complete.
SAMS/MAMS (Space & Microgravity Acceleration Measurement Systems): Ongoing.
SAMPLE: Inc16 samples have been downloaded with 15S and handed over to science team.
SLEEP (Sleep-Wake Actigraphy & Light Exposure during Spaceflight): “Garrett, per the schedule, you have completed your last week of required sleep logging. Next week you will have your last Actiwatch download/initialization session. During your last download session, the PI did not like the way the data was scaling. As of now you are not at risk of being lost as a subject, but we would like to preserve as much data as we can for later troubleshooting or ground analysis. Therefore, during this session, you will remove the Li battery in your Actiwatch to assist in ground troubleshooting. You will download, initialize, and don Peggy’s CDR Actiwatch. We greatly appreciate your participation.”
SOLAR (Solar Monitoring Observatory): The last Sun observation period has closed on 05/10. The next observation windows will start on 06/03. – SOVIM: Waiting for the Sun. During last week, the ground teams performed troubleshooting on the SOVIM covers, but with no luck; – SOLSPEC: Waiting for the Sun;
– SOLACES: Waiting for the Sun.
SPHERES (Synchronized Position Hold, Engage, Reorient, Experimental Satellite): In progress.
Swab (Characterization of Microorganisms & Allergens in Spacecraft): Complete.
TRAC (Test of Reaction & Adaptation Capabilities): Planned.
WAICO #1 (Waving and Coiling of Arabidopsis Roots at Different g-levels): The 4 EC’s of Rotor B are yet to be retrieved by the crew from the blocked Rotor B. The first troubleshooting activity, which consists of inserting back 6 Reference Experiment Containers (ECs) into the Rotor A, has been performed on 05/15. Unfortunately, while activating BIOLAB from ground after this crew activity, a caution signal was encountered with the BIOLAB Smoke Detector. Engineering teams are analyzing the telemetry. The next BIOLAB activity (Rotor B Locking Pin actuator exchange) is currently planned on 05/27.
CEO (Crew Earth Observations): Through 5/20 the ground has received a total of 1,768 of ISS/CEO images for review and cataloging. “We are currently reviewing imagery with camera times corresponding to the following target list requests: Riachao Ring Impact Crater, Brazil (not apparent yet in your views); Serra da Cangalha Impact Crater, Brazil (not apparent yet in your views); High Central Andean Glaciers (many excellent views of a number of rarely-photographed icefields and small glaciers); and Antarctic Ice Pack (poor weather and lighting reveals only a few small icebergs). Very poor lighting conditions for daylight awake passes have persisted in the Southern Hemisphere this week with your orbit tracks staying near the terminator. Hang in there! Viewing conditions and opportunities should improve dramatically next week. Your striking image of the Zion National Park area, acquired late last month, will be posted on NASA/GSFC’s Earth Observatory website this weekend. The extent and orientation of the erosion pattern and linear “joint” fractures are beautifully documented”.
CEO photography can be studied at this “Gateway” website:
http://eol.jsc.nasa.gov (as of 3/1/08, this database contained 757,605 views of the Earth from space, with 314,000 from the ISS alone).
Significant Events Ahead (all dates Eastern Time, some changes possible.):
05/31/08 — STS-124/Discovery/1J launch – JEM PM “Kibo”, racks, RMS (5:02pm EDT nominal)
06/02/08 — STS-124/Discovery/1J docking (1:49pm)
Tentative STS-124 docked working timeline:
- 6/03 — EVA-1 (11:30am, 6.5 hrs.), OBSS transfer, JPM prep, S3/S4 SARJ TBA install, JPM install
- 6/04 — JPM Vestibule outfitting, Reisman/Chamitoff handovers
- 6/05 — EVA-2 (11:30am, 6.5 hrs), JTVE install, JRMS cvr remv, NTA prep, CP 9 ETVCG retrv
- 6/06 — JLP relocate to JPM; JLP Vestibule leak check
- 6/07 — JLP Vestibule outfitting
- 6/08 — EVA-3 (10:30am, 6.3 hrs), S1 NTA install, CP9 ETVCG install, JRMS cover remove
- 6/09 — JRMS checkouts, JLP Vestibule outfitting, A/L BCM R&R
- 6/10 — Sayonara (~4:00pm), hatch close (~4:30pm)
- 6/11 — Undocking (~9:04am); OBSS survey/inspection
- 6/12 — Mostly off-duty
- 6/13 — Stowing; deorbit preps
- 6/14 — Deorbit burn (~9:56am);
06/14/08 — STS-124/Discovery landing (KSC: ~10:59am EDT, nominal)
07/10/08 — Russian EVA-20 (7/10-11)
09/05/08 — ATV1 undocking
09/09/08 — Progress M-64/29P undocking (from DC1)
09/10/08 — Progress M-65/30P launch
09/12/08 — Progress M-65/30P docking
10/01/08 — NASA 50 Years
10/08/08 — STS-125/Atlantis Hubble Space Telescope Service Mission 4 (SM4)
10/11/08 — Progress M-65/30P undocking (from SM aft port)
10/12/08 — Soyuz TMA-13/17S launch
10/14/08 — Soyuz TMA-13/17S docking (SM aft port)
10/23/08 — Soyuz TMA-12/16S undocking (FGB nadir)
11/03/08 — Soyuz TMA-13/17S relocation (from SM aft to FGB nadir)
11/10/08 — STS-126/Endeavour/ULF2 launch – MPLM Leonardo, LMC
11/20/08 — ISS 10 Years
11/26/08 — Progress M-66/31P launch
11/28/08 — Progress M-66/31P docking
12/04/08 — STS-119/Discovery/15A launch – S6 truss segment
12/06/08 — STS-119/Discovery/15A docking
12/15/08 — STS-119/Discovery/15A undocking
2QTR CY09 — STS-127/2J/A launch – JEM EF, ELM-ES, ICC-VLD
3QTR CY09 — STS-128/17A/Atlantis – MPLM(P), last crew rotation
05/??/09 — Six-person crew on ISS (following Soyuz 18S-2 docking)
3QTR CY09 — STS-129/ULF3/Discovery – ELC1, ELC2
4QTR CY09 — STS-130/20A/Endeavour – Node-3 + Cupola
1QTR CY10 — STS-131/19A/Atlantis – MPLM(P)
1QTR CY10 — STS-132/ULF4/Discovery – ICC-VLD, MRM1 (contingency)
2QTR CY10 — STS-133/ULF5/Endeavour – ELC3, ELC4 (contingency).